1. Given the linear model -9x - 7y + -1 = 0, find the residual when...

1.

Given the linear model -9x - 7y + -1 = 0, find the residual when x=0 and the actually data point is found to be (0, 7). Round to the nearest hundredth.

2.

Given the linear model 3x - 1y = -(8), the "Fit" when x=-1 would be...? Round to the nearest hundredth.

3.

Given the linear model 3x - 10y = -(-3), what value of y is predicted when x=2? Round to the nearest hundredth

4.

Given the linear model -6x + 11y = 7, the "Fit" when x=9 would be...? Round to the nearest hundredth.

5.

The following linear model is created: y = -7x+8. What is the predicted value of the dependent variable when the independent variable has a value of 2.8?
Answer #1

1)

Predicted value of y when x = 0 is -1/7

Thus, the residual = 7 - (-1/7) = 50/7 = 7.14

2)

The "Fit" when x = -1 is (3*-1 + 8) = 5

3)

The predicted value of y when x = 2 is (3*2 - 3)/10 = 0.30

4)

The "Fit" when x = 9 is (7 + 6*9)/11 = 61/11 = 5.55

5)

The predicted value of the dependent variable, y when the independent variable, x = 2.8 is (-7*2.8 + 8) = -11.6